OpenStudy (anonymous):

Integrate x^2(e^-x/2)

OpenStudy (mathmale):

Just to be certain that I'm interpreting your post correctly: Is the following what you meant?\[x ^{2e ^{\frac{ -x }{ 2 }}}?\] If not, please advise. Suggest you learn to use Equation Editor (blue button, below).

OpenStudy (anonymous):

\[x ^{2}\exp ^{ \frac{ -x }{ 2 }}\]

OpenStudy (anonymous):

Thanks hehe

OpenStudy (mathmale):

I'm going to take your word that \(x ^{2}\exp ^{ \frac{ -x }{ 2 }}\) is correct.

OpenStudy (mathmale):

And so you want \[\int\limits_{}^{}x ^{2}\exp ^{ \frac{ -x }{ 2 }}dx\] What approach or approaches would you consider using here?

OpenStudy (anonymous):

Product rule? :S

OpenStudy (mathmale):

the product rule is for differentiation, not for integration. What about integration by parts?

OpenStudy (mathmale):

Hint:\[\int\limits_{}^{}e ^{\frac{ -x }{ 2 }}dx=-2e ^{\frac{ -x }{ 2 }}+C\]

OpenStudy (anonymous):

\[u=x ^{2}\] \[u'=2x\] \[v=\exp ^{\frac{ -x }{ 2 }}\]\[v'=2\exp ^{\frac{ -x }{ 2 }}\]

OpenStudy (mathmale):

Your u and u' are fine, but it's v' that equals exp(-x/2) and it's v that equals -2exp(-x/2).

OpenStudy (mathmale):

Although I'm going to get off the Internet now, others could continue helping you with this. If no one else has helped you to your satisfaction when I'm back on OpenStudy, I'd gladly continue. Good luck!

OpenStudy (anonymous):

Okay, thanks! :)